# A Debiased Estimator for the Mediation Functional in Ultra-High-Dimensional Setting in the Presence of Interaction Effects

### v1

@@ Abstract
Mediation analysis is crucial in many fields of science for understanding the mechanisms or processes through which an independent variable affects an outcome, thereby providing deeper insights into causal relationships and improving intervention strategies. Despite advances in analyzing the mediation effect with fixed/low-dimensional mediators and covariates, our understanding of estimation and inference of mediation functional in the presence of (ultra)-high-dimensional mediators and covariates is still limited. In this paper, we present an estimator for mediation functional in a high-dimensional setting that accommodates the interaction between covariates and treatment in generating mediators, as well as interactions between both covariates and treatment and mediators and treatment in generating the response. We demonstrate that our estimator is $\sqrt{n}$-consistent and asymptotically normal, thus enabling reliable inference on direct and indirect treatment effects with asymptotically valid confidence intervals. A key technical contribution of our work is to develop a multi-step debiasing technique, which may also be valuable in other statistical settings with similar structural complexities where accurate estimation depends on debiasing. Corresponding authors.
@@ 1 Introduction
The rest of the paper is organized as follows. After introducing the problem setting in Section 2, we describe our debiased estimation procedure in Section 3. In Section 4, we provide the bias analysis of our estimator and its connection to the fixed-dimensional setting. In Section 5, we provide a theoretical justification for our approach under high-dimensional asymptotic. The proof of the main theorem, along with technical lemmas, can be found in the Appendix. We conclude in Section 6 and discuss future work.
@@ 2 Notations and Problem Setting
$Y$ $=(1-A)\left(\alpha_{0}+X^{\top}\beta_{0}+M^{\top}\gamma_{0}\right)+A\left(\alpha_{1}+X^{\top}\beta_{1}+M^{\top}\gamma_{1}\right)+\epsilon,$ (2.3)
$M$ $=(1-A)\left(\delta_{0}+\mathbf{B}_{0}X\right)+A\left(\delta_{1}+\mathbf{B}_{1}X\right)+U,$
where $\epsilon$ and $U$ are centered random shocks and the matrices $\mathbf{B}_{1},\mathbf{B}_{0}\in\mathbb{R}^{q\times p}$, and vectors $\beta_{1},\beta_{0}\in\mathbb{R}^{p}$, $\gamma_{1},\gamma_{0}\in\mathbb{R}^{q}$ contain unknown regression coefficients that encode the relationships among the exposures, mediators and outcome. Note that this data generating process (DGP) is flexible in the sense that
@@ 3 Debiased Estimator for the Mediation Functional
$\hat{\phi}_{1}=[{\beta}_{1}^{\top}\quad{\gamma}_{1}^{\top}]^{\top}=\arg\min_{\tilde{\phi}_{1}}\left\{\sum_{\{i:A_{i}=1\}}\left(Y_{i}-\mathbf{W}_{i,.}\cdot\tilde{\phi}_{1}\right)^{2}+\lambda_{1}\|\tilde{\phi}_{1}\|_{1}\right\}.$
@@ 5.1 Assumptions
The assumption $(k\vee s)\log(p+q)/\sqrt{n}=o(1)$ typically arises in debiased lasso literature (Javanmard and Montanari, , 2014; Van de Geer et al., , 2014; Zhang and Zhang, , 2014) to establish the $\sqrt{n}$-consistency of the debiased estimator. Cai and Guo, (2017) proved that this assumption is necessary unless we have more information on the design. The last assumption, i.e., $sk\log(p\vee q)/\sqrt{n_{c}\wedge n_{t}}=o(1)$, is required for the high-dimensional matrix over matrix regression (Step $1$ in our algorithm 1) and further bias correction. Whether this is necessary, we leave this as a future work.
@@
Assume that the noise is homoskedastic: $\operatorname{Var}[\epsilon_{i}^{(1)}|\mathbf{W}_{t,i}]=\sigma_{1}^{2}$ for every $i=1,\ldots,n_{t}$, and that the noise $\epsilon_{i}^{(1)}:=Y_{i}^{(1)}-\mathbb{E}[Y_{i}^{(1)}|\mathbf{W}_{t,i}]$ is i.i.d uniformly sub-Gaussian with mean 0 a finite parameter $v_{1}^{2}S_{1}>0$. Similarly, suppose that we have homoskedastic noise: $\operatorname{Var}[U_{i*}^{(0)}\mid X_{i}]=\sigma^{2}_{2}$ for all $i=1,\ldots,n_{c}$, and also the response noise $U_{i*}^{(0)}$ is i.i.d uniformly sub-Gaussian with 0 mean and a finite parameter $v_{2}^{2}S_{2}>0$.
@@
  • The operator norm $\|\mathbf{B}_{0}\|_{\text{op}}\leqslant C$, $\|\mathbf{B}_{0}\|_{1,\infty}\leq C_{B}$, $\|\mu\|_{\infty}\leqslant C_{X}$ and $\|\gamma_{1}\|_{2}\leqslant C_{\gamma}$ are finite and upper bounded by some constants $C$, $C_{B}$, $C_{X}$ and $C_{\gamma}$.

### v2

@@ Abstract
Mediation analysis is a crucial tool for uncovering the mechanisms through which a treatment affects the outcome, providing deeper causal insights and guiding effective interventions. Despite advances in analyzing the mediation effect with fixed/low-dimensional mediators and covariates, our understanding of estimation and inference of mediation functional in the presence of (ultra)-high-dimensional mediators and covariates is still limited. In this paper, we present an estimator for mediation functional in a high-dimensional setting that accommodates the interaction between covariates and treatment in generating mediators, as well as interactions between both covariates and treatment and mediators and treatment in generating the response. We demonstrate that our estimator is $\sqrt{n}$-consistent and asymptotically normal, thus enabling reliable inference on direct and indirect treatment effects with asymptotically valid confidence intervals. A key technical contribution of our work is to develop a multi-step debiasing technique, which may also be valuable in other statistical settings with similar structural complexities where accurate estimation depends on debiasing. We evaluate our proposed methodology through extensive simulation studies and apply it to the TCGA lung cancer dataset to estimate the effect of smoking, mediated by DNA methylation, on the survival time of lung cancer patients.
Corresponding authors.
@@ 1 Introduction
The rest of the paper is organized as follows. After introducing the problem setting in Section 2, we describe our debiased estimation procedure in Section 3. In Section 4, we provide the bias analysis of our estimator and its connection to the fixed-dimensional setting. In Section 5, we provide a theoretical justification for our approach under high-dimensional asymptotic. The proof of the main theorem, along with technical lemmas, can be found in the Appendix. In Section 6, we conduct a simulation experiment and a real data experiments on TCGA Lung cancer dataset. A software implementation for R is available at UHDmedi. We conclude in Section 7 and discuss future work.
@@ 2 Notations and Problem Setting
$Y$ $=(1-A)\left(\alpha_{0}+X^{\top}\beta_{0}+M^{\top}\gamma_{0}+{\epsilon}^{\prime}\right)+A\left(\alpha_{1}+X^{\top}\beta_{1}+M^{\top}\gamma_{1}+{\epsilon}\right),$ (2.3)
$M$ $=(1-A)\left(\delta_{0}+\mathbf{B}_{0}X+U\right)+A\left(\delta_{1}+\mathbf{B}_{1}X+U^{\prime}\right),$
where $(\epsilon,{\epsilon}^{\prime})$ and $(U,U^{\prime})$ are centered random shocks and the matrices $\mathbf{B}_{1},\mathbf{B}_{0}\in\mathbb{R}^{q\times p}$, and vectors $\beta_{1},\beta_{0}\in\mathbb{R}^{p}$, $\gamma_{1},\gamma_{0}\in\mathbb{R}^{q}$ contain unknown regression coefficients that encode the relationships among the exposures, mediators and outcome. Note that this data generating process (DGP) is flexible in the sense that
@@ 3 Debiased Estimator for the Mediation Functional
$\hat{\phi}_{1}=[{\hat{\beta}}_{1}^{\top}\quad{\hat{\gamma}}_{1}^{\top}]^{\top}=\arg\min_{\tilde{\phi}_{1}}\left\{\sum_{\{i:A_{i}=1\}}\left(Y_{i}-\mathbf{W}_{i,.}\cdot\tilde{\phi}_{1}\right)^{2}+\lambda_{1}\|\tilde{\phi}_{1}\|_{1}\right\}.$
@@ 5.1 Assumptions
The assumption $(k\vee s)\log(p+q)/\sqrt{n}=o(1)$ typically arises in debiased lasso literature (Javanmard and Montanari, , 2014; Van de Geer et al., , 2014; Zhang and Zhang, , 2014) to establish the $\sqrt{n}$-consistency of the debiased estimator. Cai and Guo, (2017) proved that this assumption is necessary unless we have more information on the design. The last assumption, i.e., $sk\log(p\vee q)/\sqrt{n_{c}\wedge n_{t}}=o(1)$, is required for the high-dimensional matrix over matrix regression (Step $1$ in our algorithm 1) and further bias correction. Whether this is necessary, we leave this as future work.
@@
Assume that the noise is homoskedastic: $\operatorname{Var}[\epsilon_{i}\ |\ \mathbf{W}_{t,i}]=\sigma_{1}^{2}$ for every $i=1,\ldots,n_{t}$, and that the noise $\epsilon_{i}:=Y_{i}^{(1)}-\mathbb{E}[Y_{i}^{(1)}|\mathbf{W}_{t,i}]$ is i.i.d uniformly sub-Gaussian with mean 0 a finite parameter $v_{1}^{2}S_{1}>0$. Similarly, suppose that we have homoskedastic noise: $\operatorname{Var}[U_{i}\mid X_{i}]=\sigma^{2}_{2}$ for all $i=1,\ldots,n_{c}$, and also the response noise $U_{i}$ is i.i.d uniformly sub-Gaussian with 0 mean and a finite parameter $v_{2}^{2}S_{2}>0$.
@@
  • The operator norm $\|\mathbf{B}_{0}\|_{\text{op}}\leqslant C$, $\|\mathbf{B}_{0}\|_{1,\infty}\leq C_{B}$, $\|\mu\|_{\infty}\leqslant C_{X}$ and $\|\gamma_{1}\|_{2}\leqslant C_{\gamma}$ are finite and upper bounded by some constants $C$, $C_{B}$, $C_{X}$ and $C_{\gamma}$, and $\|\hat{\gamma}_{1}-\gamma_{1}\|_{2}=\sqrt{k\log(q)/n_{t}}$ with high probability.
@@ 5.2 Consistency and Asymptotic Normality of $\hat{\theta}_{0}$
6 Simulation and Real Data Analysis
6.1 Simulation Designs
In this section, we present simulation studies to assess the performance of our proposed methodology in Section 3. Specifically, we compare our debiased estimator (Equation (5.3)) to a naive (non-debiased) estimator (defined at the beginning of Section 3) in terms of their root mean squared error (RMSE) and standard deviation. We consider the following data-generating processes:
  1. 1.
    We generate covariate $X\sim\mathcal{N}(\mu,\Sigma_{X})$, $\mu=0.5\mathbf{1}$. The covariance matrix is generated as $\Sigma_{X}=R\Lambda R^{\top}$, where $R$ is an orthonormal matrix and $1\leq\lambda_{min}(\Sigma_{X})\leq\lambda_{max}(\Sigma_{X})\leq 2$.
  2. 2.
    We generate treatment $A\sim\text{logit}(X^{\top}\alpha)$, where $\alpha\in\mathbb{R}^{p}$ is a sparse vector with $\|\alpha\|_{0}=5$ with nonzero elements randomly sampled from $\mathcal{U}(0,2)$.
  3. 3.
    Upon generating $(A,X)$, we generate the mediators from the model:
    $M=A(\mathbf{B}_{1}X+U^{\prime})+(1-A)(\mathbf{B}_{0}X+U)\,,$
    where $\mathbf{B}_{1},\mathbf{B}_{0}\in{\mathbb{R}}^{q\times p}$ are row-sparse matrices, where each row of both $\mathbf{B}_{0}$ and $\mathbf{B}_{1}$ contains $s=5$ (positions are randomly chosen) nonzero elements sampled from $\mathcal{U}(0.5,1)$. The errors are generated as $U,U^{\prime}\overset{i.i.d.}{\sim}\mathcal{N}(0,\sigma^{2}I_{q})$.
  4. 4.
    Finally, we generate the response variable from the model:
    $Y=A(X^{\top}\beta_{1}+M^{\top}\gamma_{1}+{\epsilon})+(1-A)(X^{\top}\beta_{0}+M^{\top}\gamma_{0}+{\epsilon}^{\prime})\,,$
    where $\beta_{0},\beta_{1}\in\mathbb{R}^{p}$ are sparse vectors with $k_{1}=5$ non-zero elements. The non-zero indices are chosen randomly, and the values corresponding to those indices are samples from $\mathcal{U}(0.5,1.5)$. Similarly, $\gamma_{0},\gamma_{1}\in\mathbb{R}^{q}$ are also sparse vectors with $k_{2}=5$ non-zero elements. As before, the non-zero indices are sampled randomly, and values corresponding to those indices are sampled from $\mathcal{U}(0.5,1.5)$. The errors are generated as ${\epsilon},{\epsilon}^{\prime}\overset{i.i.d.}{\sim}\mathcal{N}(0,\sigma^{2})$.
The number of samples $n$ varies over $\{250,500,750,1000,1250\}$. For each value of $n$ in this set, we vary $p=q$ over the set $\{50,400,750,1250\}$. Furthermore, the noise variance $\sigma^{2}$ is varied over the set $\{0.1,0.5,1\}$. Recall that our parameter of interest is $\theta_{0}=\bar{X}^{\top}(\beta+\mathbf{B}_{0}^{\top}\gamma_{1})$ where $\bar{X}$ is the sample average of the covariates. We keep $X$ and other parameters (e.g., $(\beta_{j},\gamma_{j},\alpha,\mathbf{B}_{j})$ for $j\in\{0,1\}$) fixed in the simulation study to keep the parameter of interest remains unchanged.
We follow Algorithm 1 to compute our estimator. The Lasso estimates in Step 1 and Step 3 are obtained using the R package glmnet, where the tuning parameter is selected by cross-validation using the lambda.1se rule from R function cv.glmnet (Friedman et al., (2010)). The debiasing weights in Step 2 and Step 4 by solving the corresponding optimization problem via R package pogs with $K_{1}=K_{2}=2.75$. This particular choice of $(K_{1},K_{2})$ is motivated by our sensitivity analysis, as elaborated in Table 4 - 7. This choice remains consistent with previous studies, e.g., Javanmard and Montanari, (2014).
Figure 2: Histogram and QQ plot of $\sqrt{n}(\hat{\theta}-\theta)$ under $n=1000,p+q=800,\sigma_{\epsilon}=0.5$.
The simulation results are summarized in Tables 1. Table 1 presents the root-mean-squared error (RMSE) and standard deviation (SD) of both the naive estimator and our proposed debiased estimator (computed using Algorithm 1), averaged over 400 Monte Carlo replications, and reports the corresponding standard deviations of the estimators across these replications. The results indicate that the debiased estimator consistently achieves a lower RMSE than the naive estimator, with the performance gap becoming increasingly prominent as the sample size as $n$ grows. Furthermore, the standard deviations of the two estimators are comparable (as evident from Table 1), suggesting that the improved RMSE of our debiased estimator is attributable to the successful bias correction. To further illustrate the performance of the estimators, we display histograms of the naive (red) and debiased (blue) estimators in the left column of Figure 2, based on 400 Monte Carlo replications. The estimators are centered around the true parameter and scaled by $\sqrt{n}$, where $n$ is the number of samples. The histogram clearly reveals that the naive estimator is substantially biased, while the debiased estimator is centered around zero as a result of successful bias correction. Additionally, the Q-Q plot of the debiased estimator, shown in the right column of Figure 2, aligns closely with the standard normal distribution, providing empirical evidence for its asymptotic normality and supporting our theoretical results.
RMSE (SD)
$s=5$ $k_{1}=k_{2}=5$ $\sigma^{2}=0.1$ $\sigma^{2}=0.5$ $\sigma^{2}=1$
$p+q$ $n$ Debiasing Naive Debiasing Naive Debiasing Naive
$100$ $250$ 0.3427 (0.1744) 0.6013 (0.0799) 0.7686 (0.3521) 1.8324 (0.2988) 0.5406 (0.4121) 1.7697 (0.3425)
$100$ $500$ 0.1021 (0.0709) 0.5254 (0.0516) 0.2569 (0.2159) 1.3490 (0.1752) 0.4262 (0.3794) 1.5741 (0.2385)
$100$ $750$ 0.0699 (0.0527) 0.4947 (0.0398) 0.2179 (0.2067) 1.1528 (0.1662) 0.3804 (0.2906) 1.3007 (0.1860)
$100$ $1000$ 0.0721 (0.0623) 0.4611 (0.0401) 0.2261 (0.1823) 1.1777 (0.1317) 0.3423 (0.3197) 1.3400 (0.1904)
$100$ $1250$ 0.0655 (0.0513) 0.3574 (0.0323) 0.1309 (0.1260) 0.8603 (0.1012) 0.2854 (0.2511) 1.2710 (0.1695)
$800$ $250$ 0.2022 (0.1538) 0.9979 (0.1613) 1.5864 (0.3329) 2.5851 (0.5604) 0.6420 (0.3056) 2.3841 (0.4128)
$800$ $500$ 0.0846 (0.0712) 0.8335 (0.0771) 0.3043 (0.2502) 1.5744 (0.2129) 0.4239 (0.2231) 1.6983 (0.2480)
$800$ $750$ 0.0989 (0.0682) 0.7680 (0.0690) 0.1586 (0.1534) 1.2404 (0.1553) 0.4282 (0.2365) 1.4553 (0.1822)
$800$ $1000$ 0.0802 (0.0724) 0.7192 (0.0603) 0.1576 (0.1416) 1.1444 (0.1439) 0.3656 (0.2024) 1.5415 (0.2026)
$800$ $1250$ 0.0564 (0.0557) 0.7300 (0.0504) 0.1379 (0.1377) 1.1339 (0.1294) 0.2459 (0.1871) 1.4200 (0.1586)
$1500$ $250$ 0.2897 (0.2021) 1.8804 (0.4664) 0.2780 (0.1859) 1.9778 (0.4542) 0.9987 (0.3472) 3.1874 (0.6424)
$1500$ $500$ 0.1639 (0.1175) 0.8545 (0.0993) 0.2474 (0.1802) 1.1352 (0.1865) 0.7018 (0.2680) 1.7272 (0.2433)
$1500$ $750$ 0.0813 (0.0796) 0.7858 (0.0783) 0.2334 (0.1326) 1.0232 (0.1435) 0.3798 (0.2487) 1.7840 (0.2349)
$1500$ $1000$ 0.0870 (0.0646) 0.7276 (0.0635) 0.1397 (0.1219) 0.9412 (0.1267) 0.2757 (0.1956) 1.5012 (0.1804)
$1500$ $1250$ 0.0692 (0.0635) 0.6697 (0.0584) 0.1253 (0.1163) 0.8814 (0.1075) 0.2436 (0.1939) 1.4856 (0.1703)
$2500$ $250$ 0.4518 (0.2448) 3.0100 (0.6146) 0.3964 (0.1788) 1.5023 (0.2777) 0.5233 (0.3316) 3.1644 (0.5719)
$2500$ $500$ 0.1096 (0.0679) 0.7782 (0.0820) 0.2751 (0.1556) 1.2072 (0.1662) 0.5395 (0.2453) 2.1845 (0.3075)
$2500$ $750$ 0.1868 (0.0948) 0.7804 (0.0701) 0.1735 (0.1266) 0.9912 (0.1318) 0.3212 (0.2701) 1.8911 (0.2386)
$2500$ $1000$ 0.0571 (0.0563) 0.6631 (0.0593) 0.1685 (0.1332) 1.0357 (0.1252) 0.2602 (0.2006) 1.9865 (0.2359)
$2500$ $1250$ 0.0484 (0.0393) 0.6325 (0.0495) 0.1108 (0.1026) 0.9615 (0.1080) 0.1988 (0.1821) 1.6966 (0.1944)
Table 1: Root-mean-square error (RMSE) with standard deviation (SD) in parentheses, for Debiasing and Naive estimators. Numbers are averaged over 400 replications with $K_{1}=K_{2}=2.75$.
6.2 Real Data Experiments
DNA methylation, an epigenetic modification characterized by the addition of a methyl group to the DNA molecule, is essential for regulating gene expression and is linked to various biological functions, and is often regarded as a potential mediator linking exposures to health outcomes (Fujii et al., , 2022). Recent studies suggest that smoking leads to widespread DNA methylation changes, which may contribute to cancer development and progression (Zhang et al., , 2016, 2023; Guo et al., , 2020). Meanwhile, certain smoking-associated methylation markers have been linked to increased mortality risk among lung cancer patients (Zhang et al., , 2016; Patel et al., , 2020). As an application of our developed methodology, we apply our method to the data from the lung cancer cohort of The Cancer Genome Atlas (TCGA) project to study the direct and indirect effects of smoking on the survival time of patients diagnosed with lung squamous cell carcinoma and lung adenocarcinoma. The data is publicly available at (https://xenabrowser.net/datapages/). The dataset includes DNA methylation data from 907 individuals, feature data for 1,299 individuals, and survival data for 1,145 individuals, among which $833$ individuals were common.
  • Our response variable, $Y$, is the logarithm of the survival time, where the survival time is measured as the number of days from diagnosis to death or the last follow-up date. We used the inter-quartile range to detect and remove the outliers in the response variable.
  • We selected $20$ clinically relevant features, $X$ (e.g., age, sex, etc.) for our study by excluding all those features which has more than $100$ missing values. See the Supplementary Materials for the complete list of the chosen features. We then imputed the remaining missing values; for numerical features (11), we used softImpute (Mazumder et al., , 2010) and for categorical variables (9), we used mice (Van Buuren and Groothuis-Oudshoorn, , 2011) imputation. Furthermore, we considered second, third-order and fourth-order polynomial features and interaction terms between categorical and numerical variables for better expressiveness. These operations finally lead to a $449$ dimensional covariate.
  • The DNA methylation data comprises 350,019 DNA methylation marker values ($M$), represented as continuous ratios ranging from 0 to 1, reflecting the intensity of methylation, where higher values indicate a greater degree of methylation. For more details, see (https://www.cancer.gov/ccg/research/genome-sequencing/tcga). We selected only those mediators with a standard deviation greater than 0.1 and with no missing values. This reduces the number of mediators to $124,233$.
After the data pre-processing, we are left with $n=738$ units, each with $449$ features $(X)$ and $124,333$ mediators ($M$). The treatment indicator ($A$) is the smoking status of the units ($A=1$ indicates a smoker). Among these samples, there were $42.6\%(315)$ are females and $57.3\%(423)$ are males. The patients’ ages range from 33 to 90 years, with a median age of 67 years. The median survival time is 603 days (1.65 years).
To estimate the direct and indirect effects (NDE and NIE) of smoking on the survival time of lung cancer patients, we fit the following high-dimensional linear model to the pre-processed data:
$M_{i}$ $=(1-A_{i})(\mathbf{B}_{0}X_{i}+U_{i})+A_{i}(\mathbf{B}_{1}X_{i}+U^{\prime}_{i}),\hskip 14.45377pti=1,\ldots,738,$
$Y_{i}$ $=(1-A_{i})(X_{i}^{\top}\beta_{0}+M_{i}^{\top}\gamma_{0}+{\epsilon}_{i}^{\prime})+A_{i}(X_{i}^{\top}\beta_{1}+M_{i}^{\top}\gamma_{1}+{\epsilon}_{i})\,.$ (6.1)
[…]
Estimate Bootstrap Mean Quantile CI (95%)
Indirect Effect -0.2486 -0.3319 (-0.7266, -0.0042)
Direct Effect -0.1022 -0.0730 (-0.5840, 0.2796)
Total Effect -0.3508 -0.4049 (-0.8297, -0.0254)
Table 2: Summary of the effects with mean and confidence interval (CI) under 300-sample bootstrap.
We compared the results with a fixed-dimensional OLS-based approach. Note that if both $X$ and $M$ are low-dimensional in Equation (6.2), we can easily estimate the NDE, NIE (and consequently the total effect) by computing ordinary least squared (OLS) esimators via regressing i) $Y$ on $(X,M)$ (on treatment and control observations separately), and ii) $M$ on $X$ (on treatment and control observations separately). To make the problem finite-dimensional, we selected CpG sites (cg19757631, cg05147638, cg24720672, cg08108679, cg05575921, cg24859433) which were identified as significant mediators in previous studies (Cui et al., , 2021; Zhang et al., 2021a, ; Zhang et al., 2021b, ), as our $M$, and $20$ clinically relevant features as our $X$ (i.e., we do not consider their higher-order terms and interactions).
In Table 3, we present the results in a fixed-dimensional setting. The results reveal that the OLS-based approach produces conclusions about the indirect and total effects that are contrary to intuition. The estimate of the indirect and total effect imply that smokers have $128.80\%$ ($e^{0.2531}\times 100\%$) (mediated through DNA methylation), and $126.52\%$ ($e^{0.2353}\times 100\%$) of survival time of non-smokers, i.e., smoking increases $\approx 29\%$ $=100\times(1-e^{0.2531})\%$ (corresponds to the NIE) and $\approx 27\%$ $=100\times(1-e^{0.2353})\%$ of the survival time of lung cancer patients. Furthermore, the confidence interval for the direct effect and total effect includes zero, suggesting a lack of statistical significance. These findings further emphasize the limitations of low-dimensional mediator selection and conventional modeling approaches, reinforcing the importance of incorporating high-dimensional methodologies for accurate mediation analysis in complex biological systems.
Estimate Bootstrap Mean Quantile CI (95%)
Indirect Effect 0.2600 0.2531 (0.1279, 0.3752)
Direct Effect -0.0223 -0.0178 (-0.2496, 0.1688)
Total Effect 0.2378 0.2353 (-0.0079, 0.4346)
Table 3: Summary of the effects in OLS (gene sites from other papers) with mean and confidence interval (CI) under 300-sample bootstrap.